1/3^(2x+6)=9^(x-4)
really need steps for this. It's been years since I've had Algebra and trying to help kiddo. Thanks

Answer by scott8148(6628)

1/3 = 3^(-1) ___ 9 = 3^2

[3^(-1)]^(2x + 6) = (3^2)^(x - 4)

3^(-2x - 6) = 3^(2x - 8)

-2x - 6 = 2x - 8 ___ 2 = 4x